# Analyzing Classroom Interaction Data Using Prompt Engineering and Network Analysis

### v1

@@ 3.3 AMEN Model for Latent Position Estimation
To obtain the latent positions that serve as mediators, we employ the AMEN model (Hoff, 2021). The AMEN model is an extension of latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks like classroom interactions. This extension is particularly relevant for our analysis of classroom interactions, which are inherently directed. The AMEN model explicitly models directed relationships by providing distinct latent positions for each student as both a sender and a receiver. The AMEN model can be expressed as:
@@
For count data, a Poisson link function is typically considered. However, a Poisson model often suffers from overdispersion, where the observed variance exceeds the theoretical mean. The sender and receiver random effects can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007). However, severe overdispersion may not be fully addressable with the sender and receiver random effects.
In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section D of the Supplementary Material. To effectively address overdispersion, we propose using a negative binomial regression model, which introduces a dispersion parameter to explicitly account for the excess variability and allows the conditional variance to vary independently of the mean. By combining the negative binomial specification with sender and receiver random effects from the AMEM model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
@@
where $\mu_{ij}$ is the mean of the negative binomial distribution where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution controls the variance of the distribution, with higher values indicating greater dispersion relative to the mean. $\alpha_{i}$ and $\beta_{j}$ are sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are sender and receiver latent feature vectors, respectively.
@@ 3.4 Estimation Methods
In the first stage, we estimate the proposed AMEN model to embed the observed student interaction network into a latent space. The prior distributions of the AMEN model parameters are specified as follows:
@@
where $\sigma_{\alpha}^{2}$, $\sigma_{\beta}^{2}$, $\sigma_{z}^{2}$, $\sigma_{w}^{2}$, and $\sigma_{r}^{2}$ represent fixed prior variance terms for each corresponding parameter, and $I_{d}$ denotes the $d\times d$ identity matrix. In this study, we set $d=2$ for interpretability and visualization purposes.
We draw the posterior samples of the AMEN model parameters and latent positions using a Metropolis-Hastings-within-Gibbs sampler. Additional details on the prior means, standard deviations, and the jumping rule for each parameter, as well as the starting points and overall convergence checks, are provided in the Section A of the Supplementary Material.
Note that the inner product between latent vectors remains invariant under transformations such as rotation or reflection. To address this, we perform Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position. This alignment ensures that the latent vectors are consistently positioned, facilitating a clearer interpretation of their relationships and enhancing the validity of subsequent analyses.
@@
In the second stage, we estimate a network mediation model. Specifically, the latent positions from the estimated AMEN model serve as mediators in the regression framework outlined in Equation (2). The network mediation model is estimated through Bayesian regression, with prior distributions specified as:
[…]
where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all regression coefficients in Equation (1), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of error term. The network mediation model is estimated using the MCMCpack package (Martin et al., 2011) in R. The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and natural indirect effect (NIE), as defined in Equation (2).
@@ 3.5.1 AMEN Model Results
Figures 6(a) and 6(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations. In these representations, each student’s position is mapped as a point in two-dimensional space. The angle between student vectors indicates their interaction similarity - smaller angles represent greater similarity in interaction patterns. Additionally, vector magnitude represents each student’s relative influence within the network structure.
Figure 2(a) reveals important bridging roles within the network structure, particularly through the positions of Gio and Isabel. Gio functions as a critical connector between two distinct student subgroups: one comprising Diego, Daniel, and Armando, and another consisting of Julian, William, and Marco. Similarly, Isabel facilitates connections between the subgroup of Melissa, Natalie, and Esther and the subgroup of Julian, William, and Marco. The latent vector visualization in Figure 6(a) captures these bridging relationships effectively, positioning Gio centrally between the connected subgroups while placing Isabel in closer proximity to the subgroup of Melissa, Natalie, and Esther.
In Figure 2(c), Isabel serves a bridging role in connecting two distinct student clusters: the Keila and Kimberly group and the Marco and Julian group. In Figure 6(b), Isabel appears closer to the Keila, Kimberly, Melissa, Natalie, and Esther groups, reflecting the higher frequency of interactions within this group compared to the Marco and Julian cluster. The visualization also captures subtle relationship dynamics, such as Esther’s unidirectional engagement with William, which influences the latent space positioning of both Isabel and William through second-order interaction effects. These examples demonstrate the advantages of latent vector analysis in visualizing complex network data. By projecting interactions into a low-dimensional space, this approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections. This methodology provides valuable insights into group dynamics, particularly for applications in educational research.
@@ 3.5.2 Network Mediation Model Results
Our investigation extends to examining how interaction patterns influence students’ mathematical achievement through network mediation analysis. This analytical approach specifically evaluates how network structures mediate the relationship between gender and mathematics learning outcomes. The network mediation analysis quantifies the three key components: the natural indirect effect (NIE), the natural direct effect (NDE), and the total effect (TE). Tables 8 and 9 present the comprehensive results of these analyses.
@@
NIE 0.7758 1.2468 -1.7182 3.3649
NDE -5.475 1.5948 -8.8699 -2.48 *
TE -4.6992 0.8321 -6.4324 -3.056 *
@@
Analysis of the EOI network, presented in Table 8, indicates that NDE is statistically significant and negative, while NIE through the network is not statistically significant. This negative NDE after controlling for proficiency of pre-test as a confounder, suggests that gender has a clear negative direct effect on post-test scores, meaning female students tend to achieve lower math scores independent of the mediating EOI network structure.
Post.M Post.SD 2.5% 97.5%
NIE 2.8653 1.3964 0.5089 5.9881 *
NDE -6.9407 1.9286 -11.0408 -3.3803 *
TE -4.0754 1.1155 -6.344 -1.7559 *
Table 9: Network Mediation Model Results for EXP Network
In contrast, for the EXP network, Table 9 reveals a different pattern. The negative NDE, which is consistent with the EOI network, shows that female students’ scores tend to be lower when considering gender independently of network effects. However, the positive NIE suggests that gender influences network structure in ways that enhance female students’ performance. These findings highlight the crucial role of network interactions, particularly within the EXP network, in mediating and potentially mitigating gender-based performance differences through collaborative learning processes.
@@ 4 Discussion
In the current study, we proposed the analysis of classroom interactions from their dialogues through an innovative combination of prompt engineering and network analysis methodologies. The prompt engineering approach enables automated classification of interaction patterns without requiring extensive pre-labeled datasets, significantly improving upon traditional manual labeling methods in both efficiency and scalability. Building on this automated classification, our network analysis framework represents students as nodes and their interactions as edges, providing a comprehensive visualization and analysis of classroom information flow and engagement patterns. Through the projection of these networks onto latent spaces, we quantified structural characteristics that provide the understanding structure of classroom interactions.
@@
The network mediation analysis provides important insights into the relationship between classroom dynamics and student achievement. Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments which traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions. This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously. Our methodological integration validates the effectiveness of combining prompt engineering with network analysis for understanding how student interactions influence learning outcomes.
@@
Our findings suggest several promising directions for future research to deepen our understanding of classroom dynamics and refine analytical methodologies. A key area for expansion involves the analysis of teacher-student interactions, which we did not address in the current study. Teachers serve as essential facilitators of learning, significantly influencing classroom dynamics and student engagement patterns. Examining these teacher-student interactions would provide valuable insights into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning. Such analysis could inform the optimization of teaching strategies and contribute to enhanced classroom effectiveness.
@@
Future research could enhance the accuracy of interaction analysis by exploring model fine-tuning with high-quality labeled datasets, extending beyond our current prompt engineering approach. Such fine-tuning could significantly improve the model’s precision and generalizability in analyzing complex educational interactions. Additionally, the systematic integration of expert knowledge and carefully curated datasets into the prompt engineering process would strengthen the methodological reliability and validity. This enhanced approach could provide more nuanced insights into classroom interaction patterns while maintaining the efficiency advantages of automated analysis.
A critical direction for future research involves developing automated prompt-generation methodologies. The current reliance on human judgment for prompt creation and selection introduces potential variability and bias into the analytical process. An automated approach that selects prompts based on their likelihood of generating desired outputs would enhance both the efficiency and consistency of the labeling process. Such automation would transform this methodology into a more scalable and reliable tool for educational research, enabling broader application across diverse educational contexts while maintaining analytical rigor.
Acknowledgement
@@
We thank the editor, associate editor, and reviewers for their constructive comments. This work was partially supported by the National Research Foundation of Korea [grant number RS-2023-00217705 and RS-2024-00333701; Basic Science Research Program awarded to IHJ] and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation). Correspondence should be addressed to Ick Hoon Jin, Department of Applied Statistics, Department of Statistics and Data Science, Yonsei University, Seoul, Republic of Korea. E-mail: ijin@yonsei.ac.kr.

### v2

@@ 3.3 AMEN Model for Latent Position Estimation
To obtain the latent positions that serve as mediators, we employ the AMEN model (Hoff, 2021). The AMEN model is an extension of the latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks such as classroom interactions. This extension is particularly relevant for our analysis of classroom interactions, which are inherently directed. The AMEN model explicitly models directed relationships by providing distinct latent positions for each student as both a sender and a receiver. The AMEN model can be expressed as
@@
For count data, a Poisson link function is typically considered. However, a Poisson model often suffers from overdispersion, where the observed variance exceeds the theoretical mean. Random effects of the sender and receiver can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007). However, severe overdispersion may not be fully addressable with the sender and receiver random effects.
In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section LABEL:supp-sec:overdispersion of the Supplementary Material. To effectively address overdispersion, we propose using a negative binomial regression model, which introduces a dispersion parameter to explicitly account for the excess variability and allows the conditional variance to vary independently of the mean. By combining the negative binomial specification with sender and receiver random effects from the AMEN model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
@@
where $\mu_{ij}$ is the mean of the negative binomial distribution, where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution control the variance of the distribution, with higher values indicating a greater dispersion relative to the mean. $\alpha_{i}$ and $\beta_{j}$ are random effects of the sender and the receiver, respectively, and $z_{i}$ and $w_{j}$ are latent vectors of the characteristics of the sender and the receiver, respectively.
@@ 3.4 Estimation Methods
In the first stage, we estimate the proposed AMEN model to embed the observed student-interaction network into a latent space. The prior distributions of the AMEN model parameters are specified as follows:
@@
where $\sigma_{\alpha}^{2}$, $\sigma_{\beta}^{2}$, $\sigma_{z}^{2}$, $\sigma_{w}^{2}$, and $\sigma_{r}^{2}$ represent fixed prior variance terms for each corresponding parameter, and $I_{d}$ denotes the $d\times d$ identity matrix. We applied a data-driven approach to determine the latent space dimension $d$ for the AMEN model by conducting a systematic comparison of the model across candidate values $d\in\{2,3,4,5\}$ with the Bayesian Information Criterion (BIC), Deviance Information Criterion (DIC), and Watanabe-Akaike Information Criterion (WAIC) (see Table LABEL:supp-tab:dimension_selection in Section LABEL:supp-sec:amen of the Supplementary Material). We selected $d=2$ for the EXP (Explain Own Idea) network and $d=3$ for the EOI (Engage with Others’ Ideas) network, balancing goodness-of-fit and model complexity. Note that dimension selection in latent space modeling is an evolving area of research with no standard rules, and we therefore employed multiple information criteria to compare model performance across different dimensions.
We draw posterior samples of the AMEN model parameters and latent positions using MCMC. Specifically, the Metropolis-adjusted Langevin algorithm (MALA) is applied to update the latent positions $z$ and $w$, while the remaining parameters are updated using a Metropolis-Hastings-within-Gibbs sampler. Additional details on the prior means, standard deviations, and the jumping rules for the model parameters, as well as the starting points and overall convergence checks, are provided in Section LABEL:supp-sec:amen of the Supplementary Material.
Since the inner product between latent vectors remains invariant under transformations such as rotation or reflection, we performed Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position. This alignment ensures that the latent vectors are consistently positioned, facilitating a clearer interpretation of their relationships and enhancing the validity of subsequent analyses.
@@
In the second stage, we estimate a network mediation model. Specifically, the latent positions of the estimated AMEN model serve as mediators in the regression framework outlined in Equation (5). The network mediation model is estimated through Bayesian regression, with prior distributions specified as:
$\Theta\sim N(\bm{0},I),\quad\epsilon_{i}\sim N(0,\sigma^{2}),\quad\sigma^{2}\sim\text{Inverse-gamma}(0.001,0.001),$
where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all the regression coefficients in Equation (4), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of the error term. The network mediation model is estimated using the MCMCpack package (Martin et al., 2011) in R. The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and the natural indirect effect (NIE), as defined in Equation (5).
@@ 3.5.1 AMEN Model Results
Figures 5(a) and 5(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations. In these representations, each student’s position is mapped as a point in two-dimensional space. The angle between the student vectors indicates their interaction similarity; smaller angles represent greater similarity in interaction patterns. Additionally, the magnitude of the vector represents each student’s relative influence within the network structure.
The results suggest that Gio takes a particularly significant role as the primary intermediary linking two distinct clusters: the first comprising Diego and Armando, and the second encompassing Julian, William, and Marco. Similarly, Isabel performs a complementary bridging function by connecting the cluster of Keila, Kimberly and Esther with the Julian-William-Marco subgroup. The latent vector visualization presented in Figure 5(a) provides empirical support for these structural relationships, positioning Gio at a central location equidistant from both connected subgroups, while Isabel appears in closer proximity to the Keila-Kimberly-Esther cluster, indicating her primary affiliation while maintaining her bridging capacity to the adjacent subgroup.
Figure 3(c) shows that Isabel serves as a bridge in connecting two distinct student groups: the Keila and Kimberly group and the Gio and Julian group. In Figure 5(b), Isabel is positioned between the Gio–William–Julian group and the Kimberly–Esther–Natalie cluster, indicating her intermediate role in the network. Daniel also serves as a bridge, linking the Kevin and Julian group and the Maria and Jessica group. These examples demonstrate the advantages of latent vector analysis in visualizing complex network data. By projecting interactions into a low-dimensional space, the present approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections. Thereby, this methodology provides useful insights into group dynamics in education settings like the current example.
@@ 3.5.2 Network Mediation Model Results
Our investigation extends to examining how interaction patterns influence students’ mathematical achievement through network mediation analysis. This analytical approach specifically evaluates how network structures mediate the relationship between gender and mathematics learning outcomes. Network mediation analysis quantifies the three key components: NIE, NDE, and TE.
To ensure robustness and reliability, we conducted the analysis using 10 independent MCMC chains and applied convergence diagnostics. In addition, to ensure consistent parameter interpretation across chains, all posterior samples were aligned with the chain exhibiting the lowest deviation information criterion (DIC) value, using the Maximum A Posteriori (MAP) estimate as the reference point for parameter alignment. This alignment procedure was completed before computing the final parameter estimates and credible intervals. Tables 10 and 11 present the comprehensive results of these analyses.
@@
NIE 1.2317 (0.0324) 0.7538 (0.0107) -0.2966 (0.0306) 2.678 (0.0432)
NDE -1.6922 (0.0319) 0.8183 (0.0107) -3.3055 (0.0442) -0.1052 (0.0308) *
TE -0.4605 (0.0098) 0.3612 (0.0046) -1.187 (0.0119) 0.2604 (0.0131)
Table 10: Network Mediation Model Results for EXP Network
Analysis of the EXP network reveals a significant pattern, as shown in Table 10. NDE achieves statistical significance with a clear negative coefficient, indicating that female students experience lower mathematics performance through direct pathways independent of network-mediated influences. Although NIE does not reach conventional statistical significance thresholds, it demonstrates a notable positive coefficient, suggesting that gender influences explanation-seeking network positioning in ways that may benefit female students’ mathematical achievement. The findings suggest that explanation networks may function as collaborative learning structures where female students can leverage peer interactions to enhance their mathematical understanding, despite facing direct performance challenges.
Post.M Post.SD 2.5% 97.5%
NIE 0.5888 (0.0662) 0.9297 (0.0162) -1.2519 (0.0864) 2.4562 (0.0497)
NDE -1.0528 (0.071) 1.1187 (0.0127) -3.1864 (0.0646) 1.1893 (0.0822)
TE -0.464 (0.0059) 0.6794 (0.0111) -1.8654 (0.0173) 0.8884 (0.032)
@@
In contrast, analysis of the EOI network presented in Table 11 demonstrates that neither NDE nor NIE through network structure achieves statistical significance, as evidenced by 95% credible intervals that encompass zero. These results suggest that within the context of the EOI network, gender differences in mathematical achievement operate through alternative pathways that are not accounted for in our current network mediation framework. The results suggest that EOI interactions do not function as a significant mediating mechanism for gender-based performance differences.
On the other hand, analysis of the EXP network reveals a contrasting pattern, as shown in Table 10. NDE achieves statistical significance with a clear negative coefficient, indicating that female students experience lower mathematics performance through direct pathways independent of network-mediated influences. Although NIE does not reach conventional statistical significance thresholds, it demonstrates a notable positive coefficient, suggesting that gender influences explanation-seeking network positioning in ways that may benefit female students’ mathematical achievement. The findings suggest that explanation networks may function as collaborative learning structures where female students can leverage peer interactions to enhance their mathematical understanding, despite facing direct performance challenges.
@@ 4 Discussion
In the current study, we proposed the analysis of classroom interactions from their dialogues through an innovative combination of prompt engineering and network analysis methodologies. The prompt engineering approach enables automated classification of interaction patterns without requiring extensive pre-labeled datasets, significantly improving upon traditional manual labeling methods in both efficiency and scalability. Our extensive validation study supports these advantages, demonstrating that LLMs achieve high inter-model consistency even on complex dialogue classification tasks where human annotators show low agreement. Despite the inherent complexity of this task, our ensemble approach produces consistent labeling results while systematically identifying uncertain utterances through Shannon entropy-based uncertainty quantification. Building on this automated classification, our network analysis framework represents students as nodes and their interactions as edges, providing a comprehensive visualization and analysis of classroom information flow and engagement patterns. Through the projection of these networks onto latent spaces, we quantified the structural characteristics that provide the structure of the understanding of classroom interactions. Our contribution extends beyond traditional approaches through the integration of ensemble-based prompt engineering with network mediation analysis. To our knowledge, this combined approach with systematic uncertainty quantification and network-based structural analysis has not been applied in educational dialogue analysis, representing a novel methodological advancement that addresses both classification reliability challenges and the limitations of utterance-level analysis through relational pattern analysis.
@@
Network mediation analysis provides important information on the relationship between classroom dynamics and student achievement. Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments that traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions. This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture the temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously. Our methodological integration demonstrates the potential of combining prompt engineering with network analysis to understand how student interactions are associated with learning outcomes, with our mediation analysis being exploratory in nature and aiming to identify patterns and associations rather than establish causal relationships.
@@
Our findings suggest several promising directions for future research to deepen our understanding of classroom dynamics and refine analytical methodologies. A key area for expansion involves the analysis of teacher-student interactions, which we did not address in the current study. Teachers serve as essential facilitators of learning, significantly influencing classroom dynamics and student engagement patterns. Examining these teacher-student interactions would provide valuable insight into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning. Such an analysis could inform the optimization of teaching strategies and contribute to increased classroom effectiveness.
An important avenue for future research involves exploring temporal variations in classroom interaction networks. Time-varying network approaches could effectively capture the evolution of interaction patterns throughout the instructional period, revealing how classroom dynamics change and develop over time. Such temporal analysis would illuminate whether the mediating effects of classroom interactions vary across different phases of the learning process, potentially yielding more nuanced insights into the dynamic nature of peer collaboration in the classroom and its impact on academic outcomes.
Our current analysis faces methodological constraints that limit causal interpretation. Specifically, our mediation model incorporates only a limited set of confounding variables, relying primarily on pre-test performance as the primary confounding adjustment. This methodological limitation likely violates the “no omitted confounders” assumption essential for robust causal mediation interpretation. The collection of more comprehensive demographic, socioeconomic, and contextual variables would enable stronger causal inference and substantially improve the validity of our mediation findings.
In addition, our current approach employs a two-stage estimation procedure due to inherent identifiability constraints within AMEN model specifications. Specifically, the rotation and reflection non-identifiability properties that characterize latent space models prevent reliable joint estimation of network formation and mediation parameters within a single unified framework. This methodological constraint introduces potential uncertainty propagation issues that may affect the precision of our mediation effect estimates.
@@
Future research should prioritize methodological enhancements that improve both the precision and generalizability of automated interaction analysis frameworks. While prompt engineering has demonstrated considerable promise, advancing beyond current approaches through model fine-tuning with high-quality labeled datasets would significantly enhance classification reliability in complex educational contexts. This advancement requires expanded human annotation efforts across diverse educational settings, systematically integrating expert pedagogical knowledge and carefully curated datasets into the model development process. Such comprehensive approaches would strengthen methodological rigor while preserving the efficiency advantages that make automated analysis attractive for large-scale educational research applications.
A particularly critical advancement involves developing automated prompt generation methodologies that reduce human dependency in the analytical pipeline. Current approaches rely heavily on human judgment for prompt creation, optimization, and selection, introducing potential variability and bias that can compromise analytical consistency across different research contexts. Automating these prompt development processes would enhance both operational efficiency and methodological consistency, transforming LLM-based classification into a more scalable and standardized tool for educational research.
Acknowledgment
@@
The authors thank the editor, associate editor, and reviewers for their constructive comments. The authors also thank Dr. Webb for providing the data and offering valuable insight for this study. This work was partially supported by the National Research Foundation of Korea [grant number NRF-2021S1A3A2A03088949, RS-2023-00217705, RS-2024-00333701; Basic Science Research Program awarded to IHJ], [grant number NRF-2025S1A5C3A0200632411; awarded to IHJ and MJ], International Joint Research Grant by Yonsei Graduate School [awarded to GK], and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation). Correspondence should be addressed to Ick Hoon Jin, Department of Applied Statistics, Department of Statistics and Data Science, Yonsei University, Seoul, Republic of Korea. E-mail: ijin@yonsei.ac.kr.